Find the equation of the curve so drawn that every point on it is equidistant from the origin and the intersection of the x-axis with the normal to the curve at the point.
1
Expert's answer
2019-12-16T09:32:17-0500

We need to Find the equation of the curve.


Let a Point on the curve as "A(x_1, y_1)"


Let Slope of the Tangent at this point = m


Slope of normal at this point is "= \frac {-1}{m}"


So, Equation of the normal is

"y - y_1 = \frac {-1}{m} (x - x_1)"

Normal intersecting the x axis, so "y = 0"


So, the above equation becomes


"- y_1 = \frac {-1}{m} (x - x_1)"

"- m y_1 = -1(x - x_1) \\
x = x_1 + my_1"

So, the point is "(x_1 + my_1, 0)"


Origin =(0,0)


Taken Point = "(x_1, y_1)"

Distance from origin to "(x_1, y_1)" = Distance from Intersection point with normal to "(x_1, y_1)"



"\sqrt {( x_1 - 0)^2 +(y_1 - 0)^2} = \sqrt {(x_1+my_1-x_1)^2+ (0- y_1)^2}"

"\sqrt {x_1^2 +y_1^2} = \sqrt {m^2 y_1^2 + y_1^2}"

Squaring on both sides


"{x_1^2 +y_1^2} = {m^2 y_1^2 + y_1^2}"


"{x_1^2 } = {m^2 y_1^2 }"

"x_1 = m y_1"

Here,

"m = \frac {dy}{dx}"

"x_1 = \frac {dy}{dx} y_1"

Transform "(x_1, y_1) \space to \space (x,y)"



"x = y \frac {dy}{dx}"


"y \space dy = x \space dx"

Integrating,


"\int y dy = \int x dx + c"

Here, c is constant



"\frac {y^2}{2} = \frac {x^2}{2} + c"


"\frac {y^2}{2} - \frac {x^2}{2} =c"

"y^2 - x^2 = 2c"

Answer:


the equation of the curve is "y^2 - x^2 = 2c" .